Problem 1.18: Blood is thicker than water Part A The density (mass divided by volume) of pure water is 1.00 g/cm that of whole blood is 1.05 g/cm and the density of seawater is 1.03 g/cm What is the mass of 5.00 L of pure water? m= Submit My Answers Give Up Part B What is the mass of 5.00 L of whole blood? m= Submit My Answers Give Up Part C What is the mass of 5.00 L of seawater? m= Submit My Answers Give Up
Answer #1

here,

volume , V = 5 L

a)

the density of pure water , pw = 1 g/cm^3 = 1000 g/l

mass , m = density * volume

m = 5 * 1000 = 5000 kg

b)

the density of whole blood , pb = 1.05 g/cm^3 = 1050 g/l

mass , m = density * volume

m = 5 * 1050 = 5250 g

c)

the density of sea water , ps = 1.03 g/cm^3 = 1030 g/l

mass , m = density * volume

m = 5 * 1030 = 5150 g